The amount of an investment
of P dollars for t years at simple interest rate r is given by
A + P = Prt.
Rewrite this formula by factoring out the greatest
common factor on the right-hand side.
b) Find A if $8300 is invested for 3 years at a simpleinterest rate of 15%.

Answer by checkley77(12844)

A+P=PRT
PRT=8,300*.25*3=$3,735 INTEREST
8,300+3,735=$12,035 IS THE TOTAL AMOUNT OF THE INVESTMENT AFTER 3 YEARS.